Marlene Bellamy purchased 400 shares of Writeline Communications stock at $ 55.36 per share using the prevailing minimum initial margin requirement of 54%. She held the stock for exactly 7 months and sold it without any brokerage costs at the end of that period. During the 7​-month holding​ period, the stock paid $ 1.64 per share in cash dividends. Marlene was charged 7.8 % annual interest on the margin loan. The minimum maintenance margin was 25%.

a. Calculate the initial value of the​ transaction, the debit balance​, and the equity position on​ Marlene's transaction.

b. For each of the following share​ prices, calculate the actual margin​ percentage, and indicate whether​ Marlene's margin account would have excess​ equity, would be​restricted, or would be subject to a margin​ call:

​(1)$44.24​, ​(2)$70.56​, and​ (3)$34.77.

c. Calculate the dollar amount of​ (1) dividends received and​ (2) interest paid on the margin loan during the 7​-month holding period.

d. Use each of the following sale prices at the end of the 7​-month holding period to calculate​ Marlene's annualized rate of return on the Writeline Communications stock​ transaction:

​(1)$50.64​, ​(2)$60.39​, and​ (3)$70.94.
Answer #1

a) The Initial Value of transaction:

Particulars Amount
No fo Shares Purchased 400
Value of each share 55.36
Total Value of the trade 22144
Initial Margin 54%
Value of Equity 11957.76
Value of Loan 10186.24
Total Value of Trade 22144

b) Margin Call = (1- Initial Margin/ 1- Maintenance Margin )

= 0.46/0.75 * 55.36

= 33.95. The trader will ask the investor to insert equity if the share price falls below $33.95.

Particulars Share Price
44.24 70.56 34.77
No of Shares 400 400 400
Total Value of the trade 17696 28224 13908
Value of the Loan 10186.24 10186.24 10186.24
Value of Equity 7509.76 18037.76 3721.76
Status Restricted Excess Equity Restricted

c) Dollar Value of the Dividends and the interest Paid

Particulars Amount
No of Shares 400
Dividend each share 1.64
Total Value 656
Interest Rate 7.8 p.a
Total value of the loan 10186.24
Interest Amount 463.5

d) Annualized rate of return

Formula: = {(Initial Investment + Gains - Losses)/Initial Investment} ^ 7/12 - 1

Initial Investment + gains - Losses = Total Value of the equity

For share price at 50.64 = (10262.26/11957.76) ^ 7/12 - 1 = -8.533

For share price at 60.39 = (14162.26/11957.76) ^ 7/12 - 1 = 10.37

For Share price at 70.94 = 18382.26/11957.76) ^ 7/12 - 1 = 28.51

Particulars Share Price
50.64 60.39 70.94
Dividends received 1.64 1.64 1.64
No of Shares 400 400 400
Total Value 20912 24812 29032
Less Interest paid 463.5 463.5 463.5
Less Principal Paid 10186.24 10186.24 10186.24
Total Value of the Equity 10262.26 14162.26 18382.26
Initial Margin 11957.76 11957.76 11957.76
Annualized return -8.533 10.37 28.51